# Sparse Feature Circuits: Discovering and Editing Interpretable Causal Graphs in Language Models

### v2

@@ 3.3 Case study: Subject–verb agreement across a relative clause
We find that inspecting small feature circuits produced by our technique can provide insights into how Pythia-70M arrives at observed behaviors. To illustrate this, we present a case study of a small feature circuit for subject–verb agreement across a relative clause (RC).
@@
Figure 4: Summary of the circuit for agreement across RC (full circuit in App. F.1). The model detects the number of the subject. Then, it detects the start of a PP/RC modifying the subject. Verb form discriminators promote particular verb inflections (singular or plural). Squares show number of feature nodes in the group and triangles show number of SAE error nodes, with the shading indicating the sum of $\hat{\text{IE}}$ terms across nodes in the group.
@@
To keep the number of nodes we need to annotate manageable, we set a relatively high node threshold of 0.1, resulting in a circuit with $86$ nodes and faithfulness .21.21.21.21 (computed as in §3.2). We summarize this circuit in Figure 4; the full circuit (as well as small circuits for other subject-verb agreement tasks) can be found in App. F.1. We depict SAE features with rectangles and SAE errors with triangles. We generally noticed circuits discovered with qualitatively better SAEs attributed a smaller proportion of the total effect to SAE error nodes; this suggests that our circuit discovery technique could be adapted into a measure of SAE quality.
@@
Our circuit depicts an interpretable algorithm wherein Pythia-70M selects appropriate verb forms via two pathways. The first pathway consists of MLP and embedding features which detect the number of the main subject and then generically promote matching verb forms. The second pathway begins the same, but moves the relevant number information to the end of the relative clause by using PP/RC boundary detectors.
@@
We find significant overlap between this circuit and the circuit we discovered for agreement across a prepositional phrase, with Pythia-70M handling these syntactically distinct structures in a mostly uniform way. In accordance with Finlayson et al. (2021), we find less overlap with our circuits for simple agreement and within RC agreement (Appendix F.1).
@@ 4 Application: Removing unintended signals from a classifier without disambiguating labels
NN classifiers often rely on unintended signals—e.g., spurious features. Nearly all prior work on this problem relies on access to disambiguating labeled data in which unintended signals are less predictive of labels than intended ones. However, some tasks have structural properties which disallow this assumption. For example, inputs for different classes might come from different data sources (Zech et al., 2018). Additionally, some have raised concerns (Ngo et al., 2024; Casper et al., 2023) that sophisticated LMs trained with human feedback (Christiano et al., 2023) in settings with easy-to-hard domain shift (Burns et al., 2023; Hase et al., 2024) will be misaligned because, in these settings, “overseer approval” and “desirable behavior” are equally predictive of training reward labels. More fundamentally, the problem with unintended signals is that they are unintended—not they are insufficiently predictive—and we would like our methods to reflect this.
We thus propose Spurious Human-interpretable Feature Trimming (Shift), where a human changes the generalization of a classifier by editing its feature circuit. We show that Shift removes sensitivity to unintended signals without access to disambiguating labeled data, or even without knowing what the signals are ahead of time.
@@
    Ablate from $M$ features judged to be task-irrelevant to obtain a classifier $C^{\prime}$.
@@
Step 3 removes the classifier’s dependence on unintended signals we can identify, but may disrupt performance for the intended signal. Step 4 can be used to restore some performance.
Accuracy
@@
Original 61.9 87.4 24.4
CBP 83.3 60.1 67.7
Shift 88.5 54.0 76.0
Shift + retrain 93.1 52.0 89.0
Neuron skyline 75.5 73.2 41.5
Feature skyline 88.5 54.3 62.9
Oracle 93.0 49.4 91.9
@@
We train a linear classifier based on Pythia-70M using the ambiguous set; see App. E.1 for details. We apply Shift by first discovering a circuit using the zero-ablation variant described in §3.1; this circuit, shown in App.F.2, contains $67$ features. We manually interpret each feature using an interface similar to that shown in App. D; namely, we inspect contexts and tokens from The Pile (Gao et al., 2020) that maximally activate the feature, as well as tokens whose log-probabilities are most affected by ablating the feature. We judge $55$ of these features to be task-irrelevant (e.g., features that promote female-associated language in biographies of women, as in Figure 18; see App. G for more examples). Although this interpretability step uses additional unlabeled data, we emphasize that we never use additional labeled data (or even additional unlabeled inputs from the classification dataset).
@@
  • Feature skyline. Instead of relying on human judgement to evaluate whether a feature should be ablated, we zero-ablate the $55$ features from our circuit that are most causally implicated in spurious feature accuracy on the balanced set.
  • Neuron skyline. The same as the feature skyline, but mean-ablating $55$ neurons.
@@
Results. We find (Table 2) that Shift near-completely removes our classifier’s dependence on gender information, with Step 3 of Shift providing the bulk of the improvement. We further find our judgements of which features are task-relevant to be highly informative: Shift without retraining matches the feature skyline.
Moreover, Shift critically relies on the use of SAE features. When applying Shift with neurons, essentially none of the neurons are interpretable, making it difficult to tell if they ought to be ablated; see Appendix G for examples. Because of this, we abandon the Shift with neurons baseline. Even using the balanced set to select neurons for removal (the neuron skyline) fails to match the performance of Shift.
Figure 5: Example clusters and features which participate in their circuits. Features are active on tokens shaded in blue and promote tokens shaded in red. (left) An example narrow induction feature recognizes the pattern $A3\dots A$ and copies information from the $3$ token. This composes with a succession feature to implement the prediction $A3\dots A\rightarrow 4$. (right) One feature promotes “to” after words which can take infinitive objects. A separate feature activates on objects of verbs or prepositions and promotes “to” as an object complement.
@@ 5 Unsupervised Circuit Discovery at Scale
So far, we have relied on human-collected data to specify LM behaviors for analysis. However, LMs implement numerous interesting behaviors, many of which may be counterintuitive to humans. In this section, we adapt our techniques to produce a near fully-automated interpretability pipeline, starting from raw text data and ending with thousands of feature circuits for auto-discovered model behaviors.
We do this in two steps:
@@
    Behavior discovery. We implement variants of the quanta discovery approach from Michaud et al. (2023), which work by clustering contexts based on vectors derived from Pythia-70M activations, gradients or both. Implementation details can be found in App. H.
@@
    Circuit discovery. Given a cluster, we apply the zero-ablation variant of our technique from §3 using dataset $\mathcal{D}=\{(x_{i},y_{i})\}$, the set of contexts in the cluster together with the next token appearing in The Pile, and metric $m=-\log P(y_{i}|x_{i})$.
We present example clusters, as well as interesting features participating in their associated circuits (Figure 5). Full annotated circuits can be found in App. F.3, and an interface for exploring all of our clusters and (unlabeled) circuits can be found at feature-circuits.xyz.
@@
Feature circuits can also shed interesting light on their clusters. For example, while the clusters in Figure 5 seem at first to each represent a single mechanism, circuit-level analysis reveals in both cases a union of distinct mechanisms. For cluster 475, Pythia-70M determines whether “to [verb]” is an appropriate object in two distinct manners (see Figure 5 caption). And for cluster 382, the prediction of successors relies on general succesion features, as well as multiple narrow induction features which recognize patterns like “$A3\dots A$”.
@@ 6 Related Work
Feature disentanglement. In addition to the recent work SAE work of Cunningham et al. (2024); Bricken et al. (2023), other approaches to feature disentanglement include (Schmidhuber, 1992; Desjardins et al., 2012; Kim & Mnih, 2018; Chen et al., 2016; Makhzani & Frey, 2013; He et al., 2022; Peebles et al., 2020; Schneider & Vlachos, 2021; Burgess et al., 2018; Chen et al., 2018; Higgins et al., 2017, inter alia).
@@ 7 Conclusion
We have introduced a method for discovering circuits on sparse features. Using this method, we discover human-interpretable causal graphs for a subject–verb agreement task, a classifier, and thousands of general token prediction tasks where no clear right or wrong answer exists. We can edit the set of features that models have access to by ablating sparse features that humans deem spurious; we find that this is significantly more effective than a neuron-based ablation method which has an unfair advantage.
@@ Acknowledgments
We thank Stephen Casper, Buck Schlegeris, Ryan Greenblatt, and Neel Nanda for discussion of ideas upstream to the experiments in §4. We thank Logan Riggs and Jannik Brinkmann for help training SAEs. We also thank Josh Engels and Max Tegmark for discussions about clustering and sparse projections related to §5. S.M. is supported by an Open Philanthropy alignment grant. C.R. is supported by Manifund Regrants. E.J.M is supported by the NSF Graduate Research Fellowship Program (Grant No. 2141064). Y.B. is supported by the Israel Science Foundation (Grant No. 448/20) and an Azrieli Foundation Early Career Faculty Fellowship. Y.B. and D.B. are supported by a joint Open Philanthropy alignment grant. A.M. is supported by a Zuckerman postdoctoral fellowship.
Limitations
The success of our technique relies on access to SAEs for a given model. Training such SAEs currently requires a large (but one-time) upfront compute cost, which poses an obstacle to running our methods. Additionally, model components not captured by the SAEs will remain uninterpretable after applying our method.
Much of our evaluation is qualitative. While we have quantitative evidence that feature circuits are useful for improving generalization without additional data (§4), evaluating dictionaries and circuits without downstream tasks is challenging.
Appendix A Methodological Details
@@ A.1 Computing Edge Weights
Let $e$ be an edge between an upstream node ${\mathbf{u}}$ and downstream node ${\mathbf{d}}$, corresponding to an upstream activation ${\mathbf{a}}^{u}$ and downstream activation ${\mathbf{a}}^{d}$. When estimating the causal importance of $e$, we consider two cases:
  1. (a)
    $e$ connects a layer $\ell$ residual stream component and a layer $\ell$ attention or MLP component, or $e$ connects a layer $\ell$ attention or MLP component and a layer $\ell+1$ residual stream component; [^4] As Pythia models employ parallel attention, the layer $\ell$ attention components have no effect on the layer $\ell$ MLP components.
  2. (b)
    $e$ connects a layer $\ell$ residual stream component and a layer $\ell+1$ residual stream component.
In case (a), the indirect effect of $e$ corresponds to the result of intervening to set ${\mathbf{d}}={\mathbf{d}}\left(x_{\text{clean}}|\text{do}({\mathbf{u}}={\mathbf{u}}_{\text{patch}})\right)$, but not intervening on any other downstream node on which ${\mathbf{a}}^{u}$ has a direct effect. As with nodes, we use a linear approximation:
$\hat{\text{IE}}(m;e;x_{\text{clean}},x_{\text{patch}})$ $=\left.\nabla_{{\mathbf{d}}}m\right|_{{\mathbf{d}}_{\text{clean}}}\left.\nabla_{{\mathbf{u}}}{\mathbf{d}}\right|_{{\mathbf{u}}_{\text{clean}}}\left({\mathbf{u}}_{\text{patch}}-{\mathbf{u}}_{\text{clean}}\right)$ (5)
$\approx m\left(x_{\text{clean}}\middle|\text{do}\left({\mathbf{d}}={\mathbf{d}}\left(x_{\text{clean}}\middle|\text{do}\left({\mathbf{u}}={\mathbf{u}}_{\text{patch}}\right)\right)\right)\right)$
@@
Figure 6: Computing edge weights. We compute all gradients between adjacent components by backpropagating from the activation of a downstream feature activation. Then, when computing contributions to non-adjacent nodes, we correct for indirect contributions by subtracting off a product of Jacobians. Finally, we compute estimated indirect effects via specific paths by multiplying gradients w.r.t. $m$ by the source node’s activation difference. We only keep edges with an estimated effect above edge threshold $T_{E}$ (a hyperparameter).
In case (b), we would like to give $e$ a score that represents the effect that ${\mathbf{a}}^{u}$ has on $m$ via ${\mathbf{a}}^{d}$, but which is not already explained by compositions of edges of edges of type (a) via intermediate activations ${\mathbf{a}}^{m}$. As is shown in Figure 6, for each intermediate activation ${\mathbf{a}}^{m}$, we can correct for the nontransitive effect through ${\mathbf{a}}^{m}$ by subtracting off a product of Jacobians. In general, we set
$\hat{\text{IE}}(m;e;x_{\text{clean}},x_{\text{patch}})=\left.\nabla_{{\mathbf{d}}}m\right|_{{\mathbf{d}}_{\text{clean}}}\left(\left.\nabla_{{\mathbf{u}}}{\mathbf{d}}\right|_{{\mathbf{u}}_{\text{clean}}}-\sum_{{\mathbf{m}}}\left.\nabla_{{\mathbf{m}}}{\mathbf{d}}\right|_{{\mathbf{m}}_{\text{clean}}}\left.\nabla_{{\mathbf{u}}}{\mathbf{m}}\right|_{{\mathbf{u}}_{\text{clean}}}\right)\left({\mathbf{u}}_{\text{patch}}-{\mathbf{u}}_{\text{clean}}\right)$
where the sum is taken over intermediate nodes ${\mathbf{m}}$ between ${\mathbf{u}}$ and ${\mathbf{d}}$.

### v3

@@ 3.3 Case study: Subject-verb agreement across a relative clause
We find that inspecting small feature circuits produced by our technique can provide insights into how Pythia-70M and Gemma-2-2B arrive at observed behaviors. To illustrate this, we present a case study of relatively small feature circuits for subject-verb agreement across a relative clause (RC).
@@
Figure 4: Summary of Pythia’s (a) and Gemma 2’s (b) circuits for agreement across RC (full circuits in App. C.1). The models detect the number of the subject. Then, they detect the start of a PP/RC modifying the subject. Verb form discriminators promote particular verb inflections (singular or plural). Gemma 2 additionally uses separate features to track the number of the noun that heads the current noun phrase. Squares show number of feature nodes in the group and triangles show number of SAE error nodes, with the shading indicating the sum of $\hat{\text{IE}}$ terms across nodes in the group. As we cannot directly interpret the triangles, we rely on their positions or inclusion in other groups to label them. If the label is ambiguous, we leave the triangles outside the boxes.
@@
To keep the number of nodes we need to annotate manageable, we tune our node threshold to produce a small circuit with faithfulness $>0.2$. For Pythia, this results in a circuit with $86$ nodes and faithfulness $0.21$; for Gemma we study a circuit with $223$ nodes and faithfulness $0.21$. We summarize these circuits in Figure 4; the full circuits (as well as small circuits for other subject-verb agreement tasks) can be found in App. C.1. We depict SAE features with rectangles and SAE errors with triangles.
@@
Our circuits depict interpretable algorithms wherein both models of study select appropriate verb forms via two pathways. The first pathway consists of features which detect the number of the main subject and then generically promote matching verb forms. The second pathway begins the same, but moves the relevant number information to the end of the relative clause by using PP/RC boundary detectors. Gemma 2 also uses noun phrase (NP) number trackers, which detect the number of the noun that heads an NP and remain active on all tokens until the end of the NP; these promote matching verb forms at each position, but especially at the last token of an NP.
@@
We find significant overlap between this circuit and the circuit we discovered for agreement across a prepositional phrase, with Pythia-70M and Gemma-2-2B handling these syntactically distinct structures in a mostly uniform way. In accordance with Finlayson et al. (2021), we find less overlap with our circuits for simple agreement and within RC agreement (Appendix C.1).
@@ 4 Application: Removing unintended signals from a classifier without disambiguating labels
NN classifiers often rely on unintended signals—e.g., spurious features. Nearly all prior work on mitigating this problem relies on access to disambiguating labeled data in which unintended signals are less predictive of labels than intended ones. However, some tasks have structural properties which disallow this assumption. For example, inputs for different classes might come from different data sources (Zech et al., 2018). Additionally, some have raised concerns (Ngo et al., 2024; Casper et al., 2023) that sophisticated LMs trained with human feedback (Christiano et al., 2023) in settings with easy-to-hard domain shift (Burns et al., 2023; Hase et al., 2024) will be misaligned because, in these settings, “overseer approval” and “desirable behavior” are equally predictive of training reward labels. More fundamentally, the problem with unintended signals is that they are unintended—not they are insufficiently predictive—and we would like our methods to reflect this.
We thus propose Spurious Human-interpretable Feature Trimming (Shift), where a human inspects a classifier’s feature circuit and removes features which they judge to be task-irrelevant. We show that Shift removes sensitivity to unintended signals without access to disambiguating labeled data, or even without knowing what the signals are ahead of time.
@@
    Ablate from $C$ features judged to be task-irrelevant to obtain a classifier $C^{\prime}$.
@@
Step 3 removes the classifier’s dependence on unintended signals we can identify, but may disrupt the classifier’s performance for the intended signal. Step 4 can be used to restore some performance.
Pythia-70M Gemma-2-2B
@@
Original 61.9 87.4 24.4 67.7 81.9 18.2
CBP 83.3 60.1 67.7 90.2 50.1 86.7
Random 61.8 87.5 24.4 67.3 82.3 18.0
Shift 88.5 54.0 76.0 76.0 51.5 50.0
Shift + retrain 93.1 52.0 89.0 95.0 52.4 92.9
Neuron skyline 75.5 73.2 41.5 65.1 84.3 5.6
Feature skyline 88.5 54.3 62.9 80.8 53.7 56.7
Oracle 93.0 49.4 91.9 95.0 50.6 93.1
@@
We adapt Pythia-70M and Gemma-2-2B into classifiers by training linear classification heads with the ambiguous set; see App. E.1 for probe training details. We then discover feature circuits for these classifiers using the zero-ablation variant described in §3.1; the Pythia circuit contains $67$ features, and the Gemma circuit contains $46$. We manually interpret each feature using the Neuronpedia interface (Lin & Bloom, 2023), which displays maximally activating dataset exemplars on a large text corpus, as the features’ direct effects on output logits. We judge $55$ of the Pythia features and $43$ of the Gemma features to be task-irrelevant—e.g., features that promote female-associated language in biographies of women, as in Figure 19 (see App. D for more examples features). Although this interpretability step uses additional unlabeled data, we emphasize that we never use additional labeled data (or even additional unlabeled classification data).
To apply Shift, we zero-ablate these irrelevant features. Finally, we retrain the linear classification head with the ambiguous set using activations extracted from the ablated model. We evaluate all accuracies on the balanced set.
@@
  • Random feature ablations. Perform Shift, but using (the same number of) randomly selected SAE features instead of features selected by a human annotator.
  • Feature skyline. Instead of relying on human judgement to evaluate whether a feature should be ablated, we zero-ablate the $55$ (for Pythia) or $43$ (for Gemma) features from our circuit that are most causally implicated in spurious feature accuracy on the balanced set.
  • Neuron skyline. The same as the feature skyline, but mean-ablating $55$ or $43$ neurons.
@@
Results. We find (Table 2) that Shift almost completely removes the classifiers’ dependence on gender information for both models. In the case of Gemma (but not Pythia), the feature ablations damage model performance; however, this performance is restored (without reintroducing the bias) by further training on the ambiguous set. Comparing Shift without retraining to the feature skyline, we further observe that SHIFT optimally or near-optimally identifies the best features to remove.
Shift critically relies on the use of properly selected SAE features. When ablating random SAE features, we see essentially no effect on probe performance. When applying Shift with neurons, essentially none of the neurons are interpretable, making it difficult to tell if they ought to be ablated; see Appendix D for examples. Because of this, we abandon the Shift with neurons baseline. Even using the balanced set to automatically select neurons for removal (the neuron skyline) fails to match Shift’s performance, as the neurons most implicated in spurious feature classification are also useful for ground-truth classification.
Figure 5: Example clusters and features which participate in their circuits (see App. C.3 for the full circuits). Features are active on tokens shaded in blue and promote tokens shaded in red. (left) An example narrow induction feature recognizes the pattern $A3\dots A$ and copies information from the $3$ token. This composes with a succession feature to implement the prediction $A3\dots A\rightarrow 4$. (right) One feature promotes “to” after words which can take infinitive objects. A separate feature activates on objects of verbs or prepositions and promotes “to” as an object complement.
@@ 5 Unsupervised Circuit Discovery at Scale
Previous work on circuit analysis relied on human-collected datasets to specify LM behaviors (Wang et al., 2023; Conmy et al., 2023; Hanna et al., 2023). However, LMs implement numerous interesting behaviors, many of which may be counterintuitive to humans. In this section, we adapt our techniques to produce a near-fully-automated interpretability pipeline, starting from a large text corpus—here, a large subset of The Pile (Gao et al., 2020)—and ending with thousands of feature circuits for auto-discovered model behaviors. These experiments are performed with Pythia-70M.
We proceed in two steps:
@@
    Behavior discovery via clustering. We interpret our large text corpus as a dataset $\{(x_{i},y_{i})\}$ of contexts $x_{i}$ with ground-truth next tokens $y_{i}$. Following Michaud et al. (2023), we associate a vector ${\mathbf{v}}_{i}={\mathbf{v}}(x_{i},y_{i})$ to each sample and apply a clustering algorithm to $\{{\mathbf{v}}_{i}\}$; this segments our large corpus into a number of smaller subcorpora corresponding to the clusters. Although this approach is entirely unsupervised, many of the resulting subcorpora capture human-interpretable model behaviors, such as predicting the next number in a sequence (Fig 5). We experiment with a number of ways of assigning $(x_{i},y_{i})\mapsto{\mathbf{v}}_{i}$, such as using the training gradient $\nabla_{\theta}\log P_{\theta}(y_{i}|x_{i})$ as in Michaud et al. (2023) as well as approaches which leverage SAE activations or gradients. See App. G for details.
@@
    Circuit discovery. Given a subcorpus $\mathcal{D}=\{(x_{i},y_{i})\}$, we apply the zero-ablation variant of our feature circuit discovery technique from §3 using the dataset $\mathcal{D}$ and metric $m=-\log P(y_{i}|x_{i})$. Thus, to each subcorpus we associate a feature circuit.
We present example clusters, as well as interesting features participating in their associated circuits (Figure 5). An interface for exploring all of our clusters and (unlabeled) circuits can be found at feature-circuits.xyz.
@@
Feature circuits can also shed interesting light on their clusters. For example, while the clusters in Figure 5 seem at first to each represent a single mechanism, circuit-level analysis reveals in both cases a union of distinct mechanisms. For cluster 475, Pythia-70M determines whether “to [verb]” is an appropriate object in two distinct manners (see Figure 5 caption). And for cluster 382, the prediction of successors relies on general succession features, as well as multiple narrow induction features which recognize patterns like “$A3\dots A$”.
@@ 6 Related Work
Feature disentanglement. In addition to recent work on SAEs for LM interpretability (Cunningham et al., 2024; Bricken et al., 2023; Gao et al., 2024; Rajamanoharan et al., 2024a; b), other approaches to feature disentanglement include Schmidhuber (1992); Desjardins et al. (2012); Kim & Mnih (2018); Chen et al. (2016); Makhzani & Frey (2013); He et al. (2022); Peebles et al. (2020); Schneider & Vlachos (2021); Burgess et al. (2017); Chen et al. (2018); Higgins et al. (2017); i.a.
@@ 7 Conclusion
We have introduced a method for discovering circuits on sparse features. Using this method, we discover human-interpretable causal graphs for a subject-verb agreement task, a classifier, and thousands of general token prediction tasks. We can edit the set of features that models have access to by ablating sparse features that humans deem spurious; we find that this is significantly more effective than a neuron-based ablation method which has an unfair advantage.
8 Limitations
The success of our technique relies on access to SAEs for a given model. Training such SAEs currently requires a large (but one-time) upfront compute cost. Additionally, model components not captured by the SAEs will remain uninterpretable after applying our method.
Much of our evaluation is qualitative. While we have quantitative evidence that feature circuits are useful for improving generalization without additional data (§4), evaluating dictionaries and circuits without downstream tasks is challenging. Feature labeling is also a qualitative process; thus, labels may vary across annotators, and may vary depending on the task of interest.
Reproducibility
We release code, data and autoencoders at github.com/saprmarks/feature-circuits. Experimental details can be found in appendices A, E, and G. Our experiments are conducted entirely on open-weights models. The Gemma Scope SAEs are publicly available (Lieberum et al., 2024). Features for both SAE suites can be browsed on Neuronpedia (Lin & Bloom, 2023). Our clusters and associated feature circuits can be browsed at feature-circuits.xyz.
@@ Acknowledgments
We thank Stephen Casper, Buck Schlegeris, Ryan Greenblatt, and Neel Nanda for discussion of ideas upstream to the experiments in §4. We thank Logan Riggs and Jannik Brinkmann for help training SAEs. We also thank Josh Engels and Max Tegmark for discussions about clustering and sparse projections related to §5. S.M. is supported by an Open Philanthropy alignment grant. C.R. is supported by Manifund Regrants. E.J.M. is supported by the NSF Graduate Research Fellowship Program (Grant No. 2141064). Y.B. is supported by the Israel Science Foundation (Grant No. 448/20) and an Azrieli Foundation Early Career Faculty Fellowship. Y.B. and D.B. are supported by a joint Open Philanthropy alignment grant. A.M. is supported by a Zuckerman postdoctoral fellowship.
Appendix A Methodological Details for feature circuit discovery
@@ A.1 Computing Edge Weights
Let $e$ be an edge between an upstream node ${\mathbf{u}}$ and downstream node ${\mathbf{d}}$; let also $\mathcal{M}$ be the set of nodes ${\mathbf{m}}$ intermediate between ${\mathbf{u}}$ and ${\mathbf{d}}$. We define the weight of the edge $e$ to be the effect on the metric $m$ when intervening to set
${\mathbf{d}}={\mathbf{d}}\left(x_{\text{clean}}|\text{do}\left({\mathbf{u}}={\mathbf{u}}_{\text{patch}},{\mathbf{m}}={\mathbf{m}}_{\text{clean}}:{\mathbf{m}}\in\mathcal{M}\right)\right).$
Intuitively, this captures the indirect effect of ${\mathbf{u}}$ on $m$ via the direct effect on ${\mathbf{d}}$, but excluding effects on ${\mathbf{d}}$ mediated by some other intermediate node ${\mathbf{m}}$.
As with nodes, we employ a linear approximation:
$\hat{\text{IE}}(m;e;x_{\text{clean}},x_{\text{patch}})=\left.\nabla_{\mathbf{d}}m\right|_{{\mathbf{d}}_{\text{clean}}}\left.\nabla_{{\mathbf{u}},\text{stop}(\mathcal{M})}{\mathbf{d}}\right|_{{\mathbf{u}}_{\text{clean}}}\left({\mathbf{u}}_{\text{patch}}-{\mathbf{u}}_{\text{clean}}\right)$ (5)
where $\nabla_{{\mathbf{u}},\text{stop}(\mathcal{M})}{\mathbf{d}}$ denotes the gradient of ${\mathbf{d}}$ with respect to ${\mathbf{u}}$ when treating all ${\mathbf{m}}\in\mathcal{M}$ as constant. In practice, this is computed by applying stop-gradients to all intermediate nodes ${\mathbf{m}}$ during pytorch’s backwards pass.